OpenStudy (anonymous):

Help? For this circle, a 70 degree central angle cut off an arc of 8 in. What is the circumference of the circle?

OpenStudy (anonymous):

\[\theta = arc/radius\]

OpenStudy (anonymous):

Hint: angle must be in Radian

OpenStudy (anonymous):

After finding the radius Circum..... = 2 pie r

OpenStudy (anonymous):



OpenStudy (anonymous):

@Yahoo! What's radian?

OpenStudy (anonymous):

it is a kind of way to measure angle

OpenStudy (anonymous):

pie radian = 180 degree

OpenStudy (anonymous):

For the radian I got 1.2 rad.

OpenStudy (anonymous):

Then it's something like 1.2rad=8/r?

Directrix (directrix):

Think of 360 degrees as being analogous to 2*pi*r as the circumference of the circle. So, 70/360 = 8/(2 * pi* r) where r is the radius of the circle.

Directrix (directrix):

r = 144/(7*pi) To find the circumference, C = 2 * pi* r. Substitute for r and solve. C = 2 * pi* [ (144/(7*pi))]

OpenStudy (anonymous):

I got 41.1

Directrix (directrix):

41.143 inches approx is what I got. So, we agree.

OpenStudy (anonymous):

The choose was rounded. But thank you so much. I can't believe that is 6th grade math...

Directrix (directrix):

Sixth grade math varies. I don't know the context of the problem or how the sixth graders were taught to think about this problem.

OpenStudy (anonymous):

My sister said they've never seen this before in her entire year and I'm a sophomore and I've never seen it either. Being in Algebra II I thought I had seen it before. I guess everywhere is different though.

Directrix (directrix):

The algebraic approach - writing and solving an equation - usually comes in a Geometry class. At least, that is where I have seen it.